# Planar 3 and 2 Method for Increased Sensitivity to Through-Plane Thermal Properties

### v1

@@ 4 Experimental Details
The circuit design for 3 has been detailed in Ref. [1]. We use a SR 830 lock-in amplifier as both a current source and for the detection of the 2 or /3 voltage drop across the sensor. We use the reference channel of the SR830 lock-in amplifier for the input current in the heater. Once the 1 across the heater is balanced using a variable resistor, we note down the 1 voltage across the heater to calculate the 1 current across the heater, using the steady-state resistance of the heater. Finally, the 3 signal is measured using the 3rd harmonic of the difference between the voltage dropped across the heater/sensor and the variable resistor.
@@ 5 Verification of the plane 3 and 2 methods
Figure 2 (a) and 2 (b) show the planar 3 and planar 2 measurements on kapton film whose thermal properties have been previously determined Ref. [3] with the heater acting as the sensor in the planar 3 method and the heater and the sensor placed in two sides of the film in the planar 2 method. As seen, the theoretical predictions from the Feldman’s method (solid lines) with the pre-determined thermal properties match perfectly with the experimentally measured in-phase and out-of-phase voltages for both the 3 and 2 measurements. For the 2 measurements, as the sensor has DC current across it, the phase of the 2 voltage matches with the phase of the 2 temperature (and resistance) oscillations. Therefore, the in-phase voltage relates to the primary component of the temperature rise, which is sensitive to the heating magnitude and the specific heat capacity of the material. The out-of-phase voltage is related to the thermal lag between the heat source and the temperature rise, and is therefore primarily sensitive to the thermal conductivity. For the 3 measurement, the sinusoidal (1) current is used to sense the 2 temperature oscillations. Thus, it induces a 90∘ phase flip between the sensed 3 voltage and the 2 temperature oscillation. Therefore, the in-phase voltage is primarily sensitive to the thermal conductivity while the out-of-phase voltage is primarily sensitive to the specific heat capacity of the material.
@@ 6 Implementation of the planar 3 method in a silicon-polymer stack
To demonstrate the advantage of using the planar 3 scheme when layers with high thermal resistance are buried below layers with high thermal conductivity, we present the comparison of the measurement of the thermal conductivity of buried polymer layers in a stack as shown in Fig 3a. The stack consists of 3 layers of undoped silicon (with thermal conductivity of 140 W/mK and volumetric specific heat capacity of 1.65 MJ/m3K), each with a thickness of 300 m. In between the silicon layers is a layer of polymer that is 15 m thick. This is representative of the layering in a vertically stacked integrated circuit [7]. We use the thermal conductivity of the polymer as 1.3 W/mK (the thermal conductivity of a thermally conductive epoxy) and a specific heat capacity of 1.6 MJ/m3K for our computation.
Fig 3b presents the comparison of the frequency dependent sensitivity of the measurement to thermal conductivity of all the layers when using 3 with a 150 m line heater (dashed lines) and planar heater (solid lines). As seen in the figure, the sensitivity in the measurement using a line heater is primarily dominated by the thermal conductivity of the top silicon layer at all frequencies. This is because the heat has a tendency to spread across the top, conductive silicon layer. Moreover, if there is a finite size of the sample in the lateral dimension, this places a minimum frequency of the measurement below which the analytical solution cannot be applied, given by equation 1. In the example presented in Fig 3, we use the lateral width of the sample as 10 mm (typical of a 3D integrated circuit), which places a minimum frequency of 3.5 Hz. This means, when using the line heater, the maximum sensitivity one can obtain for the first buried polymer layer is 0.1, and there is practically no sensitivity for the second polymer layer.
@@ 7 Implementation of the planar 3 method in a battery stack
An example of the advantage of the planar 3 method is evident in the depth-resolved thermal metrology of batteries. The thermal interface resistance of the electrode-electrolyte interface [5, 9] as well as the thermal conductivity of the electrode layers [10] have been related to electrochemical phenomena such as electrochemical degradation and lithiation in lithium-ion and solid-state batteries. This series of works has established thermal metrology, particularly the 3 method, as a possible diagnostic tool for electrochemical changes within batteries. However, due to the presence of metallic current collectors (copper and aluminum) between the 3 sensor and the electrodes/interfaces of interest, the heat from the 3 sensor preferentially flows in the in-plane direction of the metallic current collectors, affecting the measurement sensitivity for the interface and electrode thermal properties. With the planar 3 method, however, the heat flow is restricted to the through-plane direction, maximizing the measurement sensitivity to the interfaces and electrodes. Figure 4 shows the in-phase measurement sensitivity to the thermal conductivity (or the thermal interface resistance) for the copper current collector, graphite anode, the anode-separator interface, and the cathode-separator interface respectively for a typical lithium-ion battery as a function of the measurement frequency. The measurement sensitivities from the traditional (line) 3 method are shown in dashed lines while those from the planar 3 method are shown in solid lines. As seen, the line 3 method is mostly sensitive to the thermal conductivity of the copper film, while the measurement sensitivities to the graphite electrode, as well as the interfaces, are significantly compromised. The measurement sensitivities for anode-separator and cathode-separator interfaces are less than 0.1 and therefore not within the measurement resolution (cite source with 0.1 as resolution). On the contrary, the planar 3 method is not sensitive to the conductivity of the copper film and is more sensitive to the thermal properties of the anode and interfaces of interest by a factor of 2, 3.5, and 3.8 respectively. Similarly, for a symmetric lithium metal solid-state cell with a lithium foil as the anode material and a ceramic solid-state electrolyte, the traditional 3 method is mostly sensitive to the thermal conductivity of the copper current collector and the lithium foil anode, as shown by the dashed lines in Figure 4 (b). The measurement sensitivity to the solid-solid interface between the lithium metal anode and the solid-state electrolyte is less than 0.1. However, with the implementation of the planar 3 method, the measurement sensitivity of the interface increases to 0.5, which is an improvement by a factor of 5.

### v2

@@ 4 Experimental Details
The circuit design for 3 has been detailed in Ref. [1]. We use a SR 830 lock-in amplifier as both a current source and for the detection of the 2 or 3 voltage drop across the sensor. We use the reference channel of the SR830 lock-in amplifier for the input current in the heater. Once the 1 across the heater is balanced using a variable resistor, we note down the 1 voltage across the heater to calculate the 1 current across the heater, using the steady-state resistance of the heater. Finally, the 3 signal is measured using the 3rd harmonic of the difference between the voltage dropped across the heater/sensor and the variable resistor.
@@ 5 Verification of the plane 3 and 2 methods
Figure 2 (a) and 2 (b) show the planar 3 and planar 2 measurements on kapton film whose thermal properties have been previously determined in Ref. [3] with the heater acting as the sensor in the planar 3 method and the heater and the sensor placed in two sides of the film in the planar 2 method. As seen, the theoretical predictions from the Feldman’s method (solid lines) with the pre-determined thermal properties match perfectly with the experimentally measured in-phase and out-of-phase voltages for both the 3 and 2 measurements. For the 2 measurements, as the sensor has DC current across it, the phase of the 2 voltage matches with the phase of the 2 temperature (and resistance) oscillations. Therefore, the in-phase voltage relates to the primary component of the temperature rise, which is sensitive to the heating magnitude and the specific heat capacity of the material. The out-of-phase voltage is related to the thermal lag between the heat source and the temperature rise, and is therefore primarily sensitive to the thermal conductivity. For the 3 measurement, the sinusoidal (1) current is used to sense the 2 temperature oscillations. Thus, it induces a 90∘ phase flip between the sensed 3 voltage and the 2 temperature oscillation. Therefore, the in-phase voltage is primarily sensitive to the thermal conductivity while the out-of-phase voltage is primarily sensitive to the specific heat capacity of the material.
@@ 6 Implementation of the planar 3 method in a silicon-polymer stack
To demonstrate the advantage of using the planar 3 scheme when layers with high thermal resistance are buried below layers with high thermal conductivity, we present the comparison of the measurement of the thermal conductivity of buried polymer layers in a stack as shown in Figure 3 (a). The stack consists of 3 layers of undoped silicon (with thermal conductivity of 140 W/mK and volumetric specific heat capacity of 1.65 MJ/m3K), each with a thickness of 300 m. In between the silicon layers is a layer of polymer that is 15 m thick. This is representative of the layering in a vertically stacked integrated circuit [7]. We use the thermal conductivity of the polymer as 1.3 W/mK (the thermal conductivity of a thermally conductive epoxy) and a specific heat capacity of 1.6 MJ/m3K for our computation.
Figure 3 (b) presents the comparison of the frequency dependent sensitivity of the measurement to thermal conductivity of all the layers when using 3 with a 150 m line heater (dashed lines) and planar heater (solid lines). As seen in the figure, the sensitivity in the measurement using a line heater is primarily dominated by the thermal conductivity of the top silicon layer at all frequencies. This is because the heat has a tendency to spread across the top, conductive silicon layer. Moreover, if there is a finite size of the sample in the lateral dimension, this places a minimum frequency of the measurement below which the analytical solution cannot be applied, given by equation 1. In the example presented in Figure 3, we use the lateral width of the sample as 10 mm (typical of a 3D integrated circuit), which places a minimum frequency of 13.7 Hz. This means, when using the line heater, the maximum sensitivity one can obtain for the first buried polymer layer is 0.1, and there is practically no sensitivity for the second polymer layer.
@@ 7 Implementation of the planar 3 method in a battery stack
An example of the advantage of the planar 3 method is evident in the depth-resolved thermal metrology of batteries. The thermal interface resistance of the electrode-electrolyte interface [5, 9] as well as the thermal conductivity of the electrode layers [10] have been related to electrochemical phenomena such as electrochemical degradation and lithiation in lithium-ion and solid-state batteries. This series of works has established thermal metrology, particularly the 3 method, as a possible diagnostic tool for electrochemical changes within batteries. However, due to the presence of metallic current collectors (copper and aluminum) between the 3 sensor and the electrodes/interfaces of interest, the heat from the 3 sensor preferentially flows in the in-plane direction of the metallic current collectors, affecting the measurement sensitivity for the interface and electrode thermal properties. With the planar 3 method, however, the heat flow is restricted to the through-plane direction, maximizing the measurement sensitivity to the interfaces and electrodes. Figure 4 shows the in-phase measurement sensitivity to the thermal conductivity (or the thermal interface resistance) for the copper current collector, graphite anode, the anode-separator interface, and the cathode-separator interface respectively for a typical lithium-ion battery as a function of the measurement frequency. The measurement sensitivities from the traditional (line) 3 method are shown in dashed lines while those from the planar 3 method are shown in solid lines. As seen, the line 3 method is mostly sensitive to the thermal conductivity of the copper film, while the measurement sensitivities to the graphite electrode, as well as the interfaces, are significantly compromised. The measurement sensitivities for anode-separator and cathode-separator interfaces are less than 0.1 and therefore not within the measurement resolution. On the contrary, the planar 3 method is not sensitive to the conductivity of the copper film and is more sensitive to the thermal properties of the anode and interfaces of interest by a factor of 2, 3.5, and 3.8 respectively. Similarly, for a symmetric lithium metal solid-state cell with a lithium foil as the anode material and a ceramic solid-state electrolyte, the traditional 3 method is mostly sensitive to the thermal conductivity of the copper current collector and the lithium foil anode, as shown by the dashed lines in Figure 4 (b). The measurement sensitivity to the solid-solid interface between the lithium metal anode and the solid-state electrolyte is less than 0.1. However, with the implementation of the planar 3 method, the measurement sensitivity of the interface increases to 0.5, which is an improvement by a factor of 5.